Given that u="\chi"2m and v="\chi" n2 further let u and v be independent variables obtain the probability density of F=(U/M)/(V/N) where m and n are the df for U and V


1
Expert's answer
2021-10-14T11:56:23-0400

Let us first determine the marginal distribution of random variables U and V. For convenience, let "c=m\/2" and "d=n\/2". To compute the distribution of "F", we need to find the joint distribution of "U" and "V".

Note that,

"f_U(u)=(1/(\varGamma(c)*2^c))*u^{c-1}e^{-u/2},\space u\gt0" and

"f_V(v)=(1/(\varGamma(d)*2^d))*v^{d-1}e^{-v/2},\space v\gt0"

Since "U" and "V" are independent, their joint distribution is given as,

"f_{U,V}(u,v)=(1/\varGamma(c)\varGamma(d)2^{c+d})*u^{c-1}v^{d-1}e^{-u/2}e^{-v/2},\space u,v\gt0"

We will first find the distribution of the random variable "U\/V" by using the "cdf" method given as,

"F_{U/V}(f)=P(U/V\leqslant f)=P(U\leqslant fV)."

Now,

"P(U\leqslant fV)=(1/(\varGamma(c)\varGamma(d)2^{c+d}))*\int^\infin_0(\int^{fv}_0u^{c-1}e^{-u/2}du)v^{d-1}e^{-v/2}dv"

"f_{U/V}(f)= F'_{U/V}(f)=(f^{c-1}/\varGamma(c)\varGamma(d))\int^\infin_0u^{c+d-1}e^{-((f+1)/2)v}dv"

Observe that,

"f(v)=(f+1)^{c+d}u^{c+d-1}e^{-((f+1)\/2)v}" is the "pdf"of a Gamma random variable with parameters, "\alpha=(c+d)" and "\beta=(f+1)/2".

Since "\int^\infin_0f(v)=1,"

"f_{U/V}(f)=(\varGamma(c+d)/\varGamma(c)\varGamma(d))*(f^{c-1}/(f+1)^{c+d})"

We now have that,

"f_F(f)=f_{(U\/m,V\/n)}(f)=f_{(n\/m)(U\/V)}(f)=(m\/n)f_{U\/V}(m\/n*f)"

"=\varGamma((m+n)/2)m^{m/2}n^{n/2}f^{(m/2)-1}/\varGamma(m/2)\varGamma(n/2)(n+mf)^{(m+n)/2},\space f\gt0"

Which is an "F" distribution with "m" numerator degrees of freedom and "n" denominator degrees of freedom.